OpenStudy (anonymous):

Please Help???! Describe the relationship between amperage, voltage, and power?

OpenStudy (kropot72):

power (watts) = voltage (volts) * current (amps)

OpenStudy (anonymous):

Kropot72 Is correct, Good luck!:D

OpenStudy (unklerhaukus):

if you wanted to include resistance as a variable (and assuming ohms law is valid) \[\large{\boxed{P=IV=I^2R=\frac {V^2}R}}\qquad\qquad\qquad \scriptsize{ V=IR}\qquad I=\frac{V}{R}\]